# Asymptotic error distribution for stochastic Runge–Kutta methods of strong order one

### v1

@@ 2. Preliminaries
where $\bar{A}(t,x,h;y)$ and $\widetilde{A}(t,x,h;y)$ are two measurable functions w.r.t $(t,x,h;y)\in\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{R}\times\mathbb{R}^{m}$. Then one can construct two Markov chains $(\bar{X}_{k},\mathcal{F}_{t_{k}})$ and $(\widetilde{X}_{k},\mathcal{F}_{t_{k}})$, $k=0,1,\ldots,N$:
$\bar{X}_{0}=\widetilde{X}_{0}=\xi,$
@@
Authors of [22] proposed a fundamental strong convergence theorem, [22, Theorem 2.1], for numerical methods applied to SDEs. By following the arguments of proving [22, Theorem 2.1], we give the strong error order between $\bar{X}_{k}$ and $\widetilde{X}_{k}$.
@@
This proof follows the same steps as in the proof of [22, Theorem 2.1], by replacing $X$ therein by $\widetilde{X}$ here. We note that [22, Assumption 2.1] are used to establish [22, Lemma 2.2] and the moment boundedness of $X(t_{k})$ therein. In fact, in the proof of [22, Theorem 2.1], only the one-step approximation error [22, Eq. (2.9)-(2.10)], the moment uniform boundedness of $X(t_{k})$ and $\bar{X}_{k}$ therein, and [22, Lemma 2.2] are used. These three conditions correspond to (A1), (A2) and (A3), respectively. Thus, the proof of this theorem can be completed by following the proof procedure of [22, Theorem 2.1]. ∎
@@ 3. Properties of stochastic Runge–Kutta method for multiplicative noise
where $C_{1}(p)$ depends on $p$ but independent of $T$ (cf. [16, Chapter 2.4]).
@@ 3.1. Stochastic Runge–Kutta method
In fact, (3.6) can be proved similarly to [17, Page 39, Eq. (3.36)]. Note that $\mathbf{E}|\zeta_{n,h}|^{q}\leq\mathbf{E}|\xi_{n}|^{q}\leq K(q)$ for any $q\geq 1$. By the Hölder inequality, for any $\varepsilon\in(0,\frac{\kappa}{p})$, we have that $\mathbf{E}|\xi_{n}^{m}-\zeta_{n,h}^{m}|^{p}=\mathbf{E}\big{(}|m\int_{0}^{1}(\zeta_{n,h}+\lambda(\xi_{n}-\zeta_{n,h}))^{m-1}\mathrm{d}\lambda|^{p}|\xi_{n}-\zeta_{n,h}|^{p}\big{)}\leq\big{(}\mathbf{E}|\xi_{n}-\zeta_{n,h}|^{\frac{p}{1-\varepsilon p/\kappa}}\big{)}^{1-\varepsilon p/\kappa}K(p,\varepsilon,m)\leq K(p,\varepsilon,m)h^{\kappa-p\varepsilon}$, which verifies (3.7). In addition, (3.8) is a direct result of (3.7) with $p=1$.
@@
$Y^{SRK}=Y^{SRK}_{t,y}(t+h):=y+h\sum\limits_{i=1}^{s_{0}}\alpha_{i}f(Z_{i})+\Delta\widehat{W}\sum\limits_{i=1}^{s_{0}}\beta_{i}g(Z_{i}).$ (3.10)
@@
where $F(Z)=(f(Z_{1})^{\top},\ldots,f(Z_{s_{0}})^{\top})^{\top}$, $G(Z)=(g(Z_{1})^{\top},\ldots,g(Z_{s_{0}})^{\top})^{\top}$, every component of $e\in\mathbb{R}^{s_{0}}$ equals to $1$, $I_{d}\in\mathbb{R}^{d\times d}$ is the identity matrix, and $\otimes$ denotes the Kronecker product of matrices.
@@
Proof.
Further, there is $K_{2}>0$ such that
$\;|\Phi_{y}(Z)-(e\otimes I_{d})y|\leq K(1+|Z|)(h+|\Delta\widehat{W}|)$
Since $h_{1}$ in Lemma 3.1 is independent of $y$ and $\omega$, we obtain the well-posedness of (3.9).
@@
Proof.
In addition, by Lemma 3.1, $\big{|}\mathbf{E}\big{[}h\sum_{i=1}^{s_{0}}\alpha_{i}f(Z_{i})\big{]}\big{|}\leq Kh\sum_{i=1}^{s_{0}}(1+\mathbf{E}|Z_{i}|)\leq K(1+|y|)h$. Thus, we arrive at $|\mathbf{E}(Y^{SRK}_{t,y}(t+h)-y)|\leq K(1+|y|)h$. Finally, [17, Page 102, Lemma 2.2] and the Hölder inequality yield the desired result. ∎
In the subsequent text of this section, we always assume that $h\leq h_{1}\wedge 1$ without extra statement. The first order mean-square convergence condition of (3.9) have been obtained in [2, Eq. (3.7)]. Here, we generalize it to the $\|\cdot\|_{L^{2p}(\Omega)}$-norm for $p\in[1,\frac{\kappa}{2}]$.
@@
Proof.
$g(Z_{j})=g(y)+\int_{0}^{1}\nabla g(y+\lambda(Z_{j}-y))(Z_{j}-y)\mathrm{d}\lambda=:g(y)+R^{1}_{g,j}.$ (3.12)
$f(Z_{j})=f(y)+R^{1}_{f,j}$ (3.13)
$Z_{i}=y+h\sum_{j=1}^{s_{0}}a_{ij}f(Z_{j})+\Delta\widehat{W}\sum_{j=1}^{s_{0}}b_{ij}g(Z_{j}),~{}i=1,\ldots,s_{0}.$ (3.14)
Plugging (3.12) into (3.14) yields
$Z_{i}=y+\Delta\widehat{W}\sum_{j=1}^{s_{0}}b_{ij}g(y)+R^{1}_{Z_{i}},$ (3.15)
where $R^{1}_{Z_{i}}=h\sum_{j=1}^{s_{0}}a_{ij}f(Z_{j})+\Delta\widehat{W}\sum_{j=1}^{s_{0}}b_{ij}R^{1}_{g,j}$. It follows from (3.13) and the Hölder inequality that $\|R^{1}_{Z_{i}}\|_{\mathbf{L}^{p}(\Omega)}\leq K(1+|y|)h$.
It follows from (3.15) and the Taylor formula that
$g(Z_{i})$ $=g(y)+\nabla g(y)(Z_{i}-y)+\tilde{R}^{2}_{g,i}=g(y)+\Delta\widehat{W}\sum_{j=1}^{s_{0}}b_{ij}\nabla g(y)g(y)+R^{2}_{g,i},$ (3.16)
$\|\mathcal{D}^{2}g(y+\lambda(Z_{i}-y))\|_{\otimes}\leq K(1+|y|^{\iota}),$ (3.17)
$f(Z_{i})=f(y)+\Delta\widehat{W}\sum_{j=1}^{s_{0}}b_{ij}\nabla f(y)g(y)+R^{2}_{f,i}$ (3.18)
with $\|R^{2}_{f,i}\|_{\mathbf{L}^{p}(\Omega)}\leq K(1+|y|^{\iota})h$. Plugging (3.13) and (3.16) into (3.14) gives
$Z_{i}=y+\Delta\widehat{W}\sum_{j=1}^{s_{0}}b_{ij}g(y)+h\sum_{j=1}^{s_{0}}a_{ij}f(y)+\Delta\widehat{W}^{2}\sum_{j=1}^{s_{0}}b_{ij}\sum_{k=1}^{s_{0}}b_{jk}\nabla g(y)g(y)+R^{2}_{Z_{i}},$ (3.19)
The Taylor formula gives
$\tilde{R}^{3}_{g,i}:=\frac{1}{2}\int_{0}^{1}(1-\lambda)^{2}\mathcal{D}^{3}g(y+\lambda(Z_{i}-y))(Z_{i}-y,Z_{i}-y,Z_{i}-y)\mathrm{d}\lambda$. Similar to (3.17), it holds that $\|\mathcal{D}^{3}g(y+\lambda(Z_{i}-y))\|_{\otimes}\leq K(1+|y|^{\iota})$. Then the assumptions on $g$ and Lemma 3.1 produce $\|\tilde{R}^{3}_{g,j}\|_{\mathbf{L}^{p}(\Omega)}\leq K(1+|y|^{\iota})h^{3/2}$. Plugging (3.19) into $\nabla g(y)(Z_{i}-y)$ and (3.15) into $\frac{1}{2}\mathcal{D}^{2}g(y)(Z_{i}-y,Z_{i}-y)$, we obtain
$g(Z_{i})=g(y)+\Delta\widehat{W}\sum_{j=1}^{s_{0}}b_{ij}\nabla g(y)g(y)+R^{3}_{g,i},$ (3.20)
where
$\;+\frac{1}{2}\mathcal{D}^{2}g(y)(R^{1}_{Z_{i}},R^{1}_{Z_{i}})+\tilde{R}^{3}_{g,i}.$ (3.21)
Plugging (3.18) and (3.20) into (3.10) and using $\alpha^{\top}e=\beta^{\top}e=1$, $\beta^{\top}(Be)=\frac{1}{2}$, we have
$Y^{SRK}_{t,y}(t+h)=y+\Delta\widehat{W}g(y)+\frac{1}{2}\Delta\widehat{W}^{2}\nabla g(y)g(y)+hf(y)+R^{1}_{Y^{SRK}},$ (3.22)
$|\mathbf{E}R^{1}_{Y^{SRK}}|\leq K(1+|y|^{\iota})h^{2},\quad\|R^{1}_{Y^{SRK}}\|_{\mathbf{L}^{p}(\Omega)}\leq K(1+|y|^{\iota})h^{3/2}.$ (3.23)
$\;Y_{t,y}(t+h)$
=:absent:=:= : $\;y+\Delta Wg(y)+\frac{1}{2}\nabla g(y)g(y)h+hf(y)+\int_{t}^{t+h}\int_{t}^{s}\nabla g(Y_{t,y}(r))g(Y_{t,y}(r))\mathrm{d}W(r)\mathrm{d}W(s)+\tilde{R}^{1,1}_{Y}.$
For any $\varphi\in\mathbf{F}$ and $p\geq 1$, it holds that $\|\varphi(Y_{t,y}(t+h))\|_{\mathbf{L}^{p}(\Omega)}\leq K(1+|y|^{\iota})$ owing to a similar estimate as (3.3). Further, the Burkholder–Davis–Gundy (BDG) inequality and the Hölder inequality yield $|\mathbf{E}\tilde{R}^{1,1}_{Y}|\leq K(1+|y|^{\iota})h^{2}$ and $\|\tilde{R}^{1,1}_{Y}\|_{\mathbf{L}^{p}(\Omega)}\leq K(1+|y|^{\iota})h^{3/2}$. Again by the Itô formula and $\int_{t}^{t+h}\int_{t}^{s}\mathrm{d}W(r)\mathrm{d}W(s)=\frac{1}{2}(\Delta W^{2}-h)$, we have that
$Y_{t,y}(t+h)=y+\Delta Wg(y)+\frac{1}{2}\Delta W^{2}\nabla g(y)g(y)+hf(y)+R^{1}_{Y}$ (3.24)
By (3.22) and (3.24),
$|\mathbf{E}(Y_{t,y}(t+h)-Y^{SRK}_{t,y}(t+h))|$ $\leq K(1+|y|^{\iota})h^{2},$
This verifies the condition (A1) of Theorem 2.1. Note that the condition (A3) of Theorem 2.1 holds for $Y_{t,y}(t+h)$ (see [22, Lemma 2.2]). In addition, condition (A2) of Theorem 2.1 is fulfilled owing to Lemma 3.2 and (3.3). Thus, the proof is completed as a result of Theorem 2.1. ∎

### v2

@@ 2.1. Fundamental strong convergence theorem
where $\bar{A}(t,x,h;y)$ and $\widetilde{A}(t,x,h;y)$ are two measurable functions w.r.t. $(t,x,h;y)\in\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{R}\times\mathbb{R}^{m}$. Then one can construct two Markov chains $(\bar{X}_{k},\mathcal{F}_{t_{k}})$ and $(\widetilde{X}_{k},\mathcal{F}_{t_{k}})$, $k=0,1,\ldots,N$: Let $\bar{X}_{0}=\widetilde{X}_{0}\in\mathbf{L}^{2}(\Omega,\mathcal{F},\mathbf{P};\mathbb{R}^{d})$, and for $k=0,\ldots,N-1$:
@@
Authors of [21] proposed a fundamental strong convergence theorem, [21, Theorem 2.1], for numerical methods applied to SDEs. Following the arguments of proving [21, Theorem 2.1], we give the strong error order between $\bar{X}_{k}$ and $\widetilde{X}_{k}$.
@@
This proof follows the same steps as in the proof of [21, Theorem 2.1], by replacing $X$ therein by $\widetilde{X}$ here. We note that [21, Assumption 2.1] is used to establish [21, Lemma 2.2] and the moment boundedness of $X(t_{k})$ therein. In fact, in the proof of [21, Theorem 2.1], only the one-step approximation error [21, Eq. (2.9)-(2.10)], the moment uniform boundedness of $X(t_{k})$ and $\bar{X}_{k}$ therein, and [21, Lemma 2.2] are used. These three conditions correspond to (A1), (A2) and (A3), respectively. Thus, the proof of this theorem can be completed by following the proof procedure of [21, Theorem 2.1]. ∎
2.2. Stable convergence in distribution
In this part, we introduce the notion of stable convergence in distribution and giving a theorem on convergence in distribution of solutions to a sequence of SDEs perturbed by processes stably converging in law.
Let $\{X_{n}\}_{n=1}^{\infty}$ be a sequence of random variables with values in a Polish space $E$, all defined on $(\Omega,\mathcal{F},\mathbf{P})$. Let $(\tilde{\Omega},\tilde{\mathcal{F}},\tilde{\mathbf{P}})$ be an extension of $(\Omega,\mathcal{F},\mathbf{P})$ and let $X$ be an $E$-valued random variable on the extension. We say that $X_{n}$ stably converges in law to $X$ in $E$, written $X_{n}\overset{stably}{\Longrightarrow}X$ in $E$, if
$\mathbf{E}[Zf(X_{n})]\to\tilde{\mathbf{E}}[Zf(X)]$
[…]
Proposition 2.2.
[10, Lemma 2.1] Let $Y$ be a random variable on another Polish space $F$. If $X_{n}\overset{stably}{\Longrightarrow}X$ in $E$, then $(Y,X_{n})\overset{stably}{\Longrightarrow}(Y,X)$ in $F\times E$.
On basis of Proposition 2.2, we present the following convergence in distribution of a family of solutions to systems of SDEs.
Theorem 2.3.
Let $\Phi^{N}=\{\Phi^{N}(t),t\in[0,T]\}$, $N\geq 1$ be $\{\mathcal{F}_{t}\}$-adapted $\mathbb{R}^{d_{2}}$-valued processes defined on $(\Omega,\mathcal{F},\mathbf{P})$, and $Y=\{Y(t),t\in[0,T]\}$ be an $\{\mathcal{F}_{t}\}$-adapted $\mathbb{R}^{d_{1}}$-valued process defined on $(\Omega,\mathcal{F},\mathbf{P})$ with almost sure continuous trajectories. Let $Z^{N}\in\mathbf{C}([0,T];\mathbb{R}^{d_{2}})$, $N\geq 1$, be the strong solution of
$Z^{N}(t)=\int_{0}^{t}B_{0}(Y(s))Z^{N}(s)\mathrm{d}s+\sum_{i=1}^{m}\int_{0}^{t}B_{i}(Y(s))Z^{N}(s)\mathrm{d}W_{i}(s)+\Phi^{N}(t),~t\in[0,T],$ (2.1)
where $B_{i}:\mathbb{R}^{d_{1}}\to\mathbb{R}^{d_{2}\times d_{2}}$ are bounded and $\mathcal{D}B_{i}\in\mathbf{F}$ (i.e., $\mathcal{D}B_{i}$ grow at most polynomially) for $i=0,\ldots,m$. Assume that the following conditions hold.
  • (1)
    There is $\beta>0$ such that $Y\in\mathbf{C}([0,T];\mathbf{L}^{p}(\Omega;\mathbb{R}^{d_{1}}))$ and $\|Y(t)-Y(s)\|_{\mathbf{L}^{p}(\Omega)}\leq K|t-s|^{\beta}$ for sufficiently large $p$.
  • (2)
    $\Phi^{N}\overset{stably}{\Longrightarrow}\Phi$ in $\mathbf{C}([0,T];\mathbb{R}^{d_{2}})$, where $\Phi$ is an $\{\tilde{\mathcal{F}_{t}}\}$-adapted process defined on $(\tilde{\Omega},\tilde{\mathcal{F}},\{\tilde{\mathcal{F}_{t}}\}_{t\in[0,T]},\tilde{\mathbf{P}})$, an extension of $(\Omega,\mathcal{F},\{\mathcal{F}_{t}\}_{t\in[0,T]},\mathbf{P})$.
  • (3)
    $\sup\limits_{N\geq 1}\sup\limits_{t\in[0,T]}\mathbf{E}|\Phi^{N}(t)|^{4}+\sup\limits_{t\in[0,T]}\tilde{\mathbf{E}}|\Phi(t)|^{4}<\infty$, and there is $\alpha>0$ such that $\sup\limits_{N\geq 1}\mathbf{E}|\Phi^{N}(t)-\Phi^{N}(s)|^{2}+\tilde{\mathbf{E}}|\Phi(t)-\Phi(s)|^{2}\leq K|t-s|^{\alpha}$ for any $t,s\in[0,T]$.
[…]
$Z(t)=\int_{0}^{t}B_{0}(Y(s))Z(s)\mathrm{d}s+\sum_{i=1}^{m}\int_{0}^{t}B_{i}(Y(s))Z(s)\mathrm{d}W_{i}(s)+\Phi(t),~t\in[0,T].$ (2.2)
Proof.
The main trouble lies in that $Z^{N}$ is not a continuous function w.r.t. $(Y,W_{1},\ldots,W_{m},\Phi^{N})$. Thus, we will prove the conclusion based on the approximation argument for convergence in distribution; see [6, Theorem 3.2].
It follows from the boundedness of $B_{i}$, $i=0,1,\ldots,m$, and $\sup\limits_{N\geq 1}\sup\limits_{t\in[0,T]}\mathbf{E}|\Phi^{N}(t)|^{4}<\infty$ that $\sup\limits_{N\geq 1}\sup\limits_{t\in[0,T]}\mathbf{E}|Z^{N}(t)|^{4}<\infty$, which together with $\sup\limits_{N\geq 1}\mathbf{E}|\Phi^{N}(t)-\Phi^{N}(s)|^{2}\leq K|t-s|^{\alpha}$ yields
$\sup_{N\geq 1}\mathbf{E}|Z^{N}(t)-Z^{N}(s)|^{2}\leq K|t-s|^{\min(1,\alpha)}.$ (2.3)
Let $Z^{N,n}=\{Z^{N,n}(t),t\in[0,T]\}$ be the strong solution of
$Z^{N,n}(t)=\int_{0}^{t}B_{0}(Y(\kappa_{n}(s)))Z^{N,n}(\kappa_{n}(s))\mathrm{d}s+\sum_{i=1}^{m}\int_{0}^{t}B_{i}(Y(\kappa_{n}(s)))Z^{N,n}(\kappa_{n}(s))\mathrm{d}W_{i}(s)+\Phi^{N}(t),$
[…]
$\;\mathbf{E}\|Z^{N,n}-Z^{N}\|_{\mathbf{C}([0,t])}^{2}$
$\leq$ $\;K\mathbf{E}\int_{0}^{t}|Z^{N,n}(\kappa_{n}(s))-Z^{N}(s)|^{2}\mathrm{d}s+K\sum_{i=0}^{m}\mathbf{E}\int_{0}^{t}|B_{i}(Y(\kappa_{n}(s)))-B_{i}(Y(s))|^{2}|Z^{N}(s)|^{2}\mathrm{d}s$
$\leq$ $\;K\mathbf{E}\int_{0}^{t}|Z^{N,n}(\kappa_{n}(s))-Z^{N}(\kappa_{n}(s))|^{2}\mathrm{d}s+K\mathbf{E}\int_{0}^{t}|Z^{N}(\kappa_{n}(s))-Z^{N}(s)|^{2}\mathrm{d}s$
$\;+K\sum_{i=0}^{m}\int_{0}^{t}\|B_{i}(Y(\kappa_{n}(s)))-B_{i}(Y(s))\|_{\mathbf{L}^{4}(\Omega)}^{2}\|Z^{N}(s)\|_{\mathbf{L}^{4}(\Omega)}^{2}\mathrm{d}s.$ (2.4)
Since $\mathcal{D}B_{i}\in\mathbf{F}$, $i=0,1,\ldots,m$, we have $|B_{i}(y_{1})-B_{i}(y_{2})|\leq K(1+|y_{1}|^{l}+|y_{2}|^{l})|y_{1}-y_{2}|$ for some $l>0$. Thus, the Hölder inequality and the assumption on $Y$ gives
$\sum_{i=0}^{m}\|B_{i}(Y(\kappa_{n}(s)))-B_{i}(Y(s))\|_{\mathbf{L}^{4}(\Omega)}^{2}\leq Kn^{-2\beta}.$
Plugging the above relation into (2.2) and using $\sup\limits_{N\geq 1}\sup\limits_{t\in[0,T]}\mathbf{E}|Z^{N}(t)|^{4}<\infty$ and (2.3), we arrive at
$\mathbf{E}\|Z^{N,n}-Z^{N}\|^{2}_{\mathbf{C}([0,t])}\leq K\int_{0}^{t}\mathbf{E}\|Z^{N,n}-Z^{N}\|^{2}_{\mathbf{C}([0,s])}\mathrm{d}s+Kn^{-\min(1,\alpha,2\beta)}$
with $K$ being independent of $N$. In this way, one has $\lim\limits_{n\to\infty}\sup\limits_{N\geq 1}\mathbf{E}\|Z^{N,n}-Z^{N}\|^{2}_{\mathbf{C}([0,T])}=0$ due to the Gronwall inequality. This indicates $Z^{N,n}$ and $Z^{N}$ satisfies the condition (A1) of [6, Theorem 3.2]
For any fixed $n\in\mathbb{N}^{+}$, define the mapping $\Gamma^{n}:\mathbf{C}([0,T];\mathbb{R}^{d_{1}})\times\mathbf{C}([0,T];\mathbb{R})^{\otimes m}\times\mathbf{C}([0,T];\mathbb{R}^{d_{2}})\to\mathbf{C}([0,T];\mathbb{R}^{d_{2}})$ which maps $(u,g_{1},\ldots,g_{m},h)$ to the solution of
$f(t)=\int_{0}^{t}B_{0}(u(\kappa_{n}(s)))f(\kappa_{n}(s))\mathrm{d}s+\sum_{i=1}^{m}\int_{0}^{t}B_{i}(u(\kappa_{n}(s)))f(\kappa_{n}(s))\mathrm{d}g_{i}(s)+h(t),~t\in[0,T].$
[…]
$Z^{\infty,n}(t)=\int_{0}^{t}B_{i}(Y(\kappa_{n}(s)))Z^{\infty,n}(\kappa_{n}(s))\mathrm{d}s+\sum_{i=1}^{m}\int_{0}^{t}B_{i}(Y(\kappa_{n}(s)))Z^{\infty,n}(\kappa_{n}(s))\mathrm{d}W_{i}(s)+\Phi(t),~t\in[0,T].$
It follows from the assumption on $\Phi$ that $\sup\limits_{t\in[0,T]}\tilde{\mathbf{E}}|Z(t)|^{4}<\infty$ and $\tilde{\mathbf{E}}|Z(t)-Z(s)|^{2}\leq K|t-s|^{1\wedge\alpha}$. Similarly to the estimate for $\mathbf{E}\|Z^{N,n}-Z^{N}\|^{2}_{\mathbf{C}([0,T])}$, one can show that $\tilde{\mathbf{E}}\|Z^{\infty,n}-Z\|^{2}_{\mathbf{C}([0,T])}\leq Kn^{-\min(1,\alpha,2\beta)}$. In this way, we have
[…]
Thus, the conditions (A2) and (A3) of [6, Theorem 3.2] are fulfilled. Thus, the proof is complete by applying [6, Theorem 3.2]. ∎
@@ 3. Properties of stochastic Runge–Kutta method for multiplicative noise
where $C_{1}(p)$ depends on $p$ but independent of $T$ (cf. [15, Chapter 2.4]).
@@ 3.1. Stochastic Runge–Kutta method
In fact, (3.6) can be proved similarly to [16, Page 39, Eq. (3.36)]. Note that $\mathbf{E}|\zeta_{n,h}|^{q}\leq\mathbf{E}|\xi_{n}|^{q}\leq K(q)$ for any $q\geq 1$. By the Hölder inequality, for any $\varepsilon\in(0,\frac{\kappa}{p})$, we have that $\mathbf{E}|\xi_{n}^{m}-\zeta_{n,h}^{m}|^{p}=\mathbf{E}\big{(}|m\int_{0}^{1}(\zeta_{n,h}+\lambda(\xi_{n}-\zeta_{n,h}))^{m-1}\mathrm{d}\lambda|^{p}|\xi_{n}-\zeta_{n,h}|^{p}\big{)}\leq\big{(}\mathbf{E}|\xi_{n}-\zeta_{n,h}|^{\frac{p}{1-\varepsilon p/\kappa}}\big{)}^{1-\varepsilon p/\kappa}K(p,\varepsilon,m)\leq K(p,\varepsilon,m)h^{\kappa-p\varepsilon}$, which verifies (3.7). In addition, (3.8) is a direct result of (3.7) with $p=1$.
@@
$Y^{SRK}_{t,y}(t+h):=y+h\sum\limits_{i=1}^{s_{0}}\alpha_{i}f(Z_{i})+\Delta\widehat{W}\sum\limits_{i=1}^{s_{0}}\beta_{i}g(Z_{i}).$ (3.10)
@@
where $F(Z)=(f(Z_{1})^{\top},\ldots,f(Z_{s_{0}})^{\top})^{\top}$, $G(Z)=(g(Z_{1})^{\top},\ldots,g(Z_{s_{0}})^{\top})^{\top}$, every component of $e\in\mathbb{R}^{s_{0}}$ equals to 111, $I_{d}\in\mathbb{R}^{d\times d}$ is the identity matrix, and $\otimes$ denotes the Kronecker product of matrices. The following lemma gives the solvability of (3.11).
@@
Since $h_{1}$ in Lemma 3.1 is independent of $y$ and $\omega$, we obtain the well-posedness of (3.9). Next, we present the moment boundedness of (3.9).
@@
We postpone the proof of Lemmas 3.1 and 3.2 to the appendix. In the subsequent text of this section, we always assume that $h\leq h_{1}\wedge 1$ without extra statement. The first order mean-square convergence condition of (3.9) have been obtained in [3, Eq. (3.7)]. Here, we generalize it to the $\|\cdot\|_{L^{2p}(\Omega)}$-norm for $p\in[1,\frac{\kappa}{2}]$.
@@
Theorem 3.3 can be proved based on Theorem 2.1 and (3.6)-(3.8), whose proof is postponed to the appendix.